A rectangular playground is 48 feet long and 21 feet wide. What is the distance in yards around the playground?

Answer by london maths tutor(243)

Perimeter of the playground = 2*48 + 2*21 = 96 + 42 = 138 feet
Conversion for foot to yard : 1 foot = (1/3) yards
Therefore, perimeter in yard = 138 * (1/3) = 46 yards
Answer: The distance around the playground is 46 yards.